Star Videos, Inc., produces short musical videos for sale to retail outlets. The company’s balance sheet accounts as of January 1 are given below.

Star Videos, Inc.
Balance Sheet
January 1
Assets
Cash $ 81,800
Accounts receivable 104,400
Inventories:
Raw materials (film, costumes) $ 17,200
Videos in process 57,400
Finished videos awaiting sale 85,800 160,400
Prepaid insurance 12,650
Studio and equipment (net) 588,000
Total assets $ 947,250
Liabilities and Stockholders’ Equity
Accounts payable $ 203,000
Retained earnings 744,250
Total liabilities and stockholders’ equity $ 947,250

Because the videos differ in length and in complexity of production, the company uses a job-order costing system to determine the cost of each video produced. Studio (manufacturing) overhead is charged to videos on the basis of camera-hours of activity. The company’s predetermined overhead rate for the year ($40 per camera-hour) is based on a cost formula that estimated $280,000 in manufacturing overhead for an estimated allocation base of 7,000 camera-hours. Any underapplied or overapplied overhead is closed to cost of goods sold. The following transactions were recorded for the year:

  1. Film, costumes, and similar raw materials purchased on account, $195,500.
  2. Film, costumes, and other raw materials issued to production, $209,500 (85% of this material was considered direct to the videos in production, and the other 15% was considered indirect).
  3. Utility costs incurred (on account) in the production studio, $85,800.
  4. Depreciation recorded on the studio, cameras, and other equipment, $106,000. Three-fourths of this depreciation related to actual production of the videos, and the remainder related to equipment used in marketing and administration.
  5. Advertising expense incurred (on account), $175,000.
  6. Salaries and wages paid in cash as follows:
Direct labor (actors and directors) $ 89,200
Indirect labor (carpenters to build sets, costume designers, and so forth) $ 82,500
Administrative salaries $ 112,800
  1. Prepaid insurance expired during the year, $9,950 (70% related to production of videos, and 30% related to marketing and administrative activities).
  2. Miscellaneous marketing and administrative expenses incurred (on account), $10,350.
  3. Studio (manufacturing) overhead was applied to videos in production. The company recorded 7,250 camera-hours of activity during the year.
  4. Videos that cost $566,000 to produce according to their job cost sheets were transferred to the finished videos warehouse to await sale and shipment.
  5. Sales for the year totaled $1,000,000 and were all on account.
  6. The total cost to produce the videos that were sold according to their job cost sheets was $611,070.
  7. Collections from customers during the year totaled $950,000.
  8. Payments to suppliers on account during the year, $556,000.
  9. Underapplied or overapplied overhead $__?__.

Required:

1. Prepare a transaction analysis that records all of the above transactions. Calculate the ending balances at December 31 for all balance sheet accounts.

2. Prepare a schedule of cost of goods manufactured for the year.

3. Prepare a schedule of cost of goods sold for the year.

4. Prepare an income statement for the year.
Answer #1

Part 1

Star Videos Inc.

Transaction Analysis

For the Year Ended December 31

Cash

Accounts receivable

Ra materials

Videos in process

Finished goods

Manufacturing overhead

Prepaid expenses

Studio & equipment (net)

=

Accounts payable

Retained earnings

Beginning balance @1/1

81800

104400

17200

57400

85800

12650

588000

203000

744250

(a) Raw material purchase

195500

195500

(b) Raw materials used

(209500)

178075

31425

(c)Utility costs

85800

85800

(d) Depreciation charges

79500

(106000)

(26500)

(e) Advertising

175000

(175000)

(f)Salaries & wages

(284500)

89200

82500

(112800)

(g) Prepaid insurance

6965

(9950)

(2985)

(h) Miscellaneous marketing

10350

(10350)

(i)Applied overhead

290000

(290000)

(j)Transfer completed videos to finished goods

(566000)

566000

(k)Sales

1000000

1000000

(l)Transfer finished goods to cost of goods sold

(611070)

(611070)

(m) Cash collections from customers

950000

(950000)

(n) Payment to suppliers

(556000)

(556000)

(o) Overapplied overhead

3810

3810

Ending balances @ 12/31

191300

154400

3200

48675

40730

0

2700

482000

113650

809355

209500*85% = 178075

209500*15% = 31425

106000*3/4 = 79500

106000*1/4 = 26500

9950*30% = 2985

9950*70% = 6965

280000/7000*7250 = 290000

Part 2

Star Videos Inc.

Schedule of Cost of Goods Manufactured

For the Year Ended December 31

Direct Materials:

Beginning raw materials inventory

17200

Add: purchases of raw materials

195500

Total raw materials available

212700

Deduct: ending raw materials inventory

3200

Raw materials used in production

209500

Deduct: indirect materials included in manufacturing overhead

31425

178075

Direct labor

89200

Manufacturing overhead

286190

overapplied overhead to work in process

3810

Total manufacturing costs

557275

Add: beginning work in process inventory

57400

614675

Deduct: ending work in process inventory

48675

Cost of goods manufactured

$566000

Part 3

Star Videos Inc.

Schedule of Cost of Goods Sold

For the Year Ended December 31

Beginning finished videos inventory

85800

Cost of goods manufactured

566000

Cost of goods available for sale

651800

Less: Ending finished videos inventory

40730

Unadjusted cost of goods sold

$611070

Less: Overapplied overhead

3810

Adjusted cost of goods sold

$607260

Part 4

Star Videos Inc.

Income Statement

For the Year Ended December 31

Sales

1000000

Cost of goods sold

607260

Gross margin

392740

Selling and administrative expenses

327635

Net operating income

$65105